# A solution to the Cauchy dual subnormality problem for a cyclic analytic $2$-isometry with defect operator of rank two

### v1

@@ 3. Proof of the Main Theorem
It is proved in [16, Theorem 2.4] that in the case when measure $\mu$ is supported at a single point , the Cauchy dual of the operator $M_{z}$ is subnormal. Also, as mentioned earlier, if the measure is supported at $\{1,-1\}$, the Cauchy dual of $M_{z}$ is subnormal. Thus, we set $\xi=\cos\theta+i\sin\theta$ where $\theta\neq 0,\pi$. The proof of the main theorem is an application of the following theorem from [16].
@@
[16, Theorem 6.4] For positive scalars $c_{1},c_{2},\dots,c_{k}$ and distinct points $\zeta_{1},\zeta_{2},\dots,\zeta_{k}$ on the unit circle $\mathbb{T}$, consider the positive Borel measure $\mu=\sum_{j=1}^{k}c_{j}\delta_{\zeta_{j}}$ on $\mathbb{T}$, where $\delta_{\zeta_{j}}$ denotes the Dirac delta measure supported at $\zeta_{j}$. Let $X(z)=(z,z^{2},\dots,z^{k})^{T}$ and $\{e_{j}\}_{j=1}^{k}$ denote the standard basis of $\mathbb{C}^{k}$. Then there exist $\alpha_{1},\alpha_{2},\dots,\alpha_{k}\in\mathbb{C}\setminus\overline{\mathbb{D}}$ and a $k\times k$ upper triangular matrix $P$ such that the Dirichlet-type space $D(\mu)$ coincides with the de Branges-Rovnyak space $H(B)$ with equality of norms, where $B=\left(\frac{p_{1}}{q},\frac{p_{2}}{q},\dots,\frac{p_{k}}{q}\right)$ and
$p_{j}(z)=\left\langle PX(z),e_{j}\right\rangle~,~j=1,\dots,k,~~~q(z)=\prod_{j=1}^{k}(z-\alpha_{j})$
Moreover, $\alpha_{1},\dots,\alpha_{k}$ are governed by
@@
In addition we record the following two results from [16] that help us to check the subnormality of the Cauchy dual of operator $M_{z}$ on $H(B)$.
@@
[16, Theorem 2.1] Let $B=(b_{1},\dots,b_{k})\in{\mathcal{S}}(\mathbb{C}^{k},\mathbb{C})$ be such that $B(0)=0$ where
@@
[16, Corollary 4.3] Assume the hypothesis of Theorem 3.2. If $\alpha_{r}\overline{\alpha_{t}}\notin[1,\infty)$ for every $1\leq r\neq t\leq k$, then $M_{z}^{\prime}$ is subnormal if and only if
$\sum_{j=1}^{k}p_{j}(\alpha_{r})\overline{p_{j}(\alpha_{t})}=0,~~~1\leq r\neq t\leq k$
As a first step towards the proof of the main theorem, we proceed with the following lemma. This lemma derives a special case of Riesz-Fejer theorem by choosing $n=2,\zeta_{1}=1,\zeta_{2}=\xi,c_{1}=c_{2}=1$ in (2.1). Though the existence of $\alpha_{1},\alpha_{2}$ is ensured by Riesz-Fejer theorem, the lemma provides certain properties of these $\alpha_{1},\alpha_{2}$ which are useful for further computations.
@@
    For all $z\in\mathbb{T}$, there exist a polynomial, say $f(z)$, of degree $4$ such that
    $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2}=\displaystyle\frac{\overline{\xi}}{z^{2}}f(z)$
@@
    Polynomial $f(z)$ has exactly two roots lying outside the closed unit disc.
@@
    Let $\alpha_{1}$ and $\alpha_{2}$ be those two roots of $f(z)$ lying outside the closed unit disc, then $\alpha_{1}\alpha_{2}=b~\xi$ for some $b>0$.
@@
    $\alpha_{1}+\alpha_{2}=a(1+\xi)$ where $a=\displaystyle\frac{3b}{b+1}$.
@@
    We simplify the expression $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2}$ to get
    $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2}$
    $=\displaystyle\frac{\overline{\xi}z^{4}-3(1+\overline{\xi})z^{3}+(8+2\cos\theta)z^{2}-3(1+\xi)z+\xi}{z^{2}}$
    $=\displaystyle\frac{\overline{\xi}f(z)}{z^{2}}$
    Thus,
    $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2}=\displaystyle\frac{\overline{\xi}f(z)}{z^{2}}$ (3.10)
@@
    Note that $f(z)$ is a monic polynomial with the property that if $\alpha$ is a root then $\frac{1}{\overline{\alpha}}$ is also a root. Therefore exactly two of the roots of $f(z)$ lie outside the closed unit disc.
@@
    The roots of $f(z)$ are $\alpha_{1},\alpha_{2},\displaystyle\frac{1}{\overline{\alpha_{1}}}$ and $\frac{1}{\overline{\alpha_{2}}}$.
    Therefore, we get,
    $\alpha_{1}\alpha_{2}\displaystyle\frac{1}{\overline{\alpha_{1}}}\frac{1}{\overline{\alpha_{2}}}=\xi^{2}\text{ and hence, }\displaystyle\frac{\alpha_{1}\alpha_{2}}{\overline{\alpha_{1}}\overline{\alpha_{2}}}=\frac{\xi}{\overline{\xi}}$
@@
    In fact, we prove that $b=|\alpha_{1}||\alpha_{2}|$.
@@
    Let $\beta_{1},\beta_{2}$ be the other two roots of $f(z)$.
    Thus, $\beta_{1}=\displaystyle\frac{1}{\overline{\alpha_{1}}}$ and $\beta_{2}=\displaystyle\frac{1}{\overline{\alpha_{2}}}$
    For $z\in\mathbb{T}$, we consider
    $\displaystyle\frac{(z-\beta_{1})(z-\beta_{2})}{z^{2}}$ $=\displaystyle\frac{z^{2}-(\beta_{1}+\beta_{2})z+\beta_{1}\beta_{2}}{z^{2}}$
    $=1-\displaystyle\frac{\beta_{1}+\beta_{2}}{z}+\frac{\beta_{1}\beta_{2}}{z^{2}}$
    $=\beta_{1}\beta_{2}\left[\overline{z}^{2}-\left(\displaystyle\frac{1}{\beta_{1}}+\displaystyle\frac{1}{\beta_{2}}\right)\overline{z}+\displaystyle\frac{1}{\beta_{1}\beta_{2}}\right]$
    But we have, $\frac{1}{\beta_{1}}=\overline{\alpha_{1}}$ and $\frac{1}{\beta_{2}}=\overline{\alpha_{2}}$. Thus, using equation (3.11), we get,
    $\frac{(z-\beta_{1})(z-\beta_{2})}{z^{2}}=\displaystyle\frac{\xi}{b}\overline{q(z)}$
    Hence, we get,
    $\frac{\bar{\xi}}{z^{2}}f(z)=\bar{\xi}\frac{\xi}{b}q(z)\overline{q(z)}=\frac{1}{b}\big|q(z)\big|^{2},\text{ for all }z\in\mathbb{T}$ (3.12)
@@
    We have $\alpha_{1}\alpha_{2}=b~\xi$ and $\alpha_{1}+\alpha_{2}+\displaystyle\frac{1}{\overline{\alpha_{1}}}+\frac{1}{\overline{\alpha_{2}}}=3(1+\xi)$.
    But, $\frac{1}{\overline{\alpha_{1}}}+\frac{1}{\overline{\alpha_{2}}}=\displaystyle\frac{\overline{\alpha_{1}}+\overline{\alpha_{2}}}{\overline{\alpha_{1}}~\overline{\alpha_{2}}}=\displaystyle\frac{\xi}{b}(\overline{\alpha_{1}}+\overline{\alpha_{2}})$.
    Hence, $\alpha_{1}+\alpha_{2}+\displaystyle\frac{\xi}{b}(\overline{\alpha_{1}}+\overline{\alpha_{2}})=3(1+\xi)$.
    Let $\alpha_{1}+\alpha_{2}=r+is$,
    then we get, b(r+is)+(cosθ+isinθ)(r-is)=3b(1+cosθ+isinθ).
    This gives, s=3bb+1sinθ=asinθ where $a=\displaystyle\frac{3b}{b+1}$.
@@
    It can be verified that $f(\xi\overline{\alpha_{1}})=0$. Further, $\xi\overline{\alpha_{1}}=\displaystyle\frac{1}{\overline{\alpha_{1}}}$ gives $|\alpha_{1}|=1$ which is not true. Also, $\xi\overline{\alpha_{1}}=\displaystyle\frac{1}{\overline{\alpha_{2}}}$ gives $|\alpha_{1}\alpha_{2}|=1$ which is also not true.
    Hence, it remains to prove that $\xi\overline{\alpha_{1}}\neq\alpha_{1}$.
    Suppose $\xi\overline{\alpha_{1}}=\alpha_{1}$.
    Let $\arg(\alpha_{i})=\theta_{i}$, for $i=1,2$.
    We get, $\theta-\theta_{1}=\theta_{1}$ and thus, $\theta=2\theta_{1}$.
    But, $\alpha_{1}\alpha_{2}=b\xi$ gives $\theta_{1}+\theta_{2}=\theta$. Thus, $\theta_{2}=\theta_{1}$. But, this is not true. Hence, $\xi\overline{\alpha_{1}}\neq\alpha_{1}$.
@@
    We use the relation between the roots and the coefficients of $f(z)$ to get,
    $\alpha_{1}\alpha_{2}+\alpha_{1}\frac{1}{\overline{\alpha_{1}}}+\alpha_{1}\frac{1}{\overline{\alpha_{2}}}+\alpha_{2}\frac{1}{\overline{\alpha_{1}}}+\alpha_{2}\frac{1}{\overline{\alpha_{2}}}+\frac{1}{\overline{\alpha_{1}}~\overline{\alpha_{2}}}=(8+2\cos\theta)\xi$
    which implies,
    $1+|a(1+\xi)|^{2}=8b-b^{2}+2b\cos\theta$ and hence,
    $1+\displaystyle\frac{9b^{2}}{(b+1)^{2}}\langle 1+\xi,1+\xi\rangle=8b-b^{2}+2b\cos\theta$
    This gives
    $b^{4}-2(3+\cos\theta)b^{3}+2(2+7\cos\theta)b^{2}-2(3+\cos\theta)b+1=0$
    Thus, $b$ satisfies the polynomial g(x)=x4-(6+2cosθ)x3+(4+14cosθ)x2-(6+2cosθ)x+1.
The proof of the main theorem demands different techniques based on the location of the point $\xi=\cos\theta+i\sin\theta$ on the unit circle. We therefore divide the proof into three parts. In fact, the nature of the roots $\alpha_{1},\alpha_{2}$ appearing in Lemma 3.4 when $\cos\theta=0.6$ prompts the authors to make these cases. as can be seen in the following lemma.
@@
The following statements are valid:
@@
    If $\cos\theta>0.6$ then $\arg(\alpha_{1})=\arg(\alpha_{2})$. Further, after appropriate notations for $\alpha_{1}$ and $\alpha_{2}$, there exist $k>1$ such that $\alpha_{2}=k\alpha_{1}$.
@@
    Suppose $\cos\theta<0.6$ and $\arg(\alpha_{1})=\arg(\alpha_{2})$.
    As $\alpha_{1}+\alpha_{2}=a(1+\xi)$, we must have $\alpha_{1}=s(1+\xi)$ and $\alpha_{2}=t(1+\xi)$ for some $s,t\in\mathbb{R}$.
    This gives $\alpha_{1}+\alpha_{2}=(s+t)(1+\xi)=a(1+\xi)$ and $\alpha_{1}\alpha_{2}=st(1+\xi)^{2}=b\xi$.
@@
    Let $D$ be the discriminant of polynomial $h(x)$ then we get,
    $D=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1+\cos\theta}$
    We have, $\cos\theta<0.6$, which gives $\frac{-1}{1+\cos\theta}<\frac{-1}{1.6}$.
    Thus,
    $D$ $=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1+\cos\theta}$
    $<\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1.6}$
    $=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{5b}{4}$
    $=\displaystyle\frac{36b^{2}-5b(b^{2}+2b+1)}{4(b+1)^{2}}$
    $=\displaystyle\frac{-b(5b^{2}-26b+5)}{4(b+1)^{2}}$
    $=\displaystyle\frac{-b(5b-1)(b-5)}{4(b+1)^{2}}$
    Thus, $D<\displaystyle\frac{-b(5b-1)(b-5)}{4(b+1)^{2}}$.
@@
    Recall from part (7) of Lemma 3.4 that $b$ satisfy the polynomial
@@
    Hence, there exist $\beta_{0}\in(5,6)$ such that $g(\beta_{0})=0$. By the property of the polynomial, we get, $g\left(\frac{1}{\beta_{0}}\right)=0$. (Note that $\frac{1}{\beta_{0}}\in(0,1)$). We now prove that $g$ has no root in the interval $(1,5)$. For, suppose possible, $g$ has a root in the interval $(1,5)$. Let $\beta$ be the smallest such root of $g$. So, $g(\beta)=0$.

### v2

@@ 3. Proof of the Main Theorem
It is proved in [13, Theorem 2.4] that in the case when measure $\mu$ is supported at a single point, the Cauchy dual of the operator $M_{z}$ on $D(\mu)$ is subnormal. Also, as mentioned earlier, if the measure is supported at $\{1,-1\}$, the Cauchy dual of $M_{z}$ on $D(\mu)$ is subnormal. Thus, we set $\xi=\cos\theta+i\sin\theta$ where $\theta\neq 0,\pi$. The proof of the main theorem is an application of the following theorem from [13].
@@
[13, Theorem 6.4] For positive scalars $c_{1},c_{2},\dots,c_{k}$ and distinct points $\zeta_{1},\zeta_{2},\dots,\zeta_{k}$ on the unit circle $\mathbb{T}$, consider the positive Borel measure $\mu=\sum_{j=1}^{k}c_{j}\delta_{\zeta_{j}}$ on $\mathbb{T}$, where $\delta_{\zeta_{j}}$ denotes the Dirac delta measure supported at $\zeta_{j}$. Let $X(z)=(z,z^{2},\dots,z^{k})^{T}$ and $\{e_{j}\}_{j=1}^{k}$ denote the standard basis of $\mathbb{C}^{k}$. Then there exist $\alpha_{1},\alpha_{2},\dots,\alpha_{k}\in\mathbb{C}\setminus\overline{\mathbb{D}}$ and a $k\times k$ upper triangular matrix $P$ such that the Dirichlet-type space $D(\mu)$ coincides with the de Branges-Rovnyak space $H(B)$ with equality of norms, where $B=\left(\frac{p_{1}}{q},\frac{p_{2}}{q},\dots,\frac{p_{k}}{q}\right)$ and $p_{j}(z)=\left\langle PX(z),e_{j}\right\rangle~,~j=1,\dots,k,~~~q(z)=\prod_{j=1}^{k}(z-\alpha_{j}).$ Moreover, $\alpha_{1},\dots,\alpha_{k}$ are governed by
@@
In addition, we record the following two results from [13] that are used in the sequel.
@@
[13, Theorem 2.1] Let $B=(b_{1},\dots,b_{k})\in{\mathcal{S}}(\mathbb{C}^{k},\mathbb{C})$ be such that $B(0)=0$ where
@@
[13, Corollary 4.3] Assume the hypothesis of Theorem 3.2. If $\alpha_{r}\overline{\alpha_{t}}\notin[1,\infty)$ for every $1\leq r\neq t\leq k$, then $M_{z}^{\prime}$ is subnormal if and only if $\sum_{j=1}^{k}p_{j}(\alpha_{r})\overline{p_{j}(\alpha_{t})}=0,~~~1\leq r\neq t\leq k.$
As a first step towards the proof of the main theorem, we proceed with the following lemma. The lemma invokes the Riesz-Fejer theorem and derives a special case of equation (2.1) by choosing $n=2,\zeta_{1}=1,\zeta_{2}=\xi$ and $c_{1}=c_{2}=1.$ Further in this case, we make some observations which turn out to be decisive in the proof of the main theorem.
@@
Let $\xi=\cos\theta+i\sin\theta\in\mathbb{T}~(\xi\neq\pm 1)$ be a point of the unit circle $\mathbb{T}.$
@@
    There exist a polynomial, say $f(z)$, of degree $4$ such that for all $z\in\mathbb{T},$
    $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2}=\displaystyle\frac{\overline{\xi}}{z^{2}}f(z).$
@@
    The polynomial $f(z)$ has exactly two roots lying outside the closed unit disc.
@@
    Let $\alpha_{1}$ and $\alpha_{2}$ be those two roots of $f(z)$ lying outside the closed unit disc. Then $\alpha_{1}\alpha_{2}=b~\xi$ for some $b>0$.
@@
    The sum $\alpha_{1}+\alpha_{2}=a(1+\xi)$ where $a=\frac{3b}{b+1}$.
@@
    After simplifying the expression $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2},$ we get,
    (3.10) $|z-1|^{2}|z-\xi|^{2}+|z-1|^{2}+|z-\xi|^{2}=\displaystyle\frac{\overline{\xi}f(z)}{z^{2}}$
@@
    We know that $f(z)$ cannot have a root on the unit circle. Further, note that $f(z)$ is a monic polynomial with the property that if $\alpha$ is a root then $\frac{1}{\overline{\alpha}}$ is also a root. Therefore, exactly two of the roots of $f(z)$ lie outside the closed unit disc.
@@
    The roots of $f(z)$ are $\alpha_{1},\alpha_{2},\frac{1}{\overline{\alpha_{1}}}$ and $\frac{1}{\overline{\alpha_{2}}}$. Therefore, we get,
    $\alpha_{1}\alpha_{2}\displaystyle\frac{1}{\overline{\alpha_{1}}}\frac{1}{\overline{\alpha_{2}}}=\xi^{2}\text{ and hence, }\displaystyle\frac{\alpha_{1}\alpha_{2}}{\overline{\alpha_{1}}\overline{\alpha_{2}}}=\frac{\xi}{\overline{\xi}}.$
@@
    In fact, we prove that $b>0$ and so that $b=|\alpha_{1}||\alpha_{2}|$.
@@
    Let, $\beta_{1}=\frac{1}{\overline{\alpha_{1}}}$ and $\beta_{2}=\frac{1}{\overline{\alpha_{2}}}.$ For $z\in\mathbb{T}$, we have
    $\displaystyle\frac{(z-\beta_{1})(z-\beta_{2})}{z^{2}}$ $=\beta_{1}\beta_{2}\left[\overline{z}^{2}-\left(\displaystyle\frac{1}{\beta_{1}}+\displaystyle\frac{1}{\beta_{2}}\right)\overline{z}+\displaystyle\frac{1}{\beta_{1}\beta_{2}}\right].$
    But we have, $\frac{1}{\beta_{1}}=\overline{\alpha_{1}}$ and $\frac{1}{\beta_{2}}=\overline{\alpha_{2}}$. Thus, using equation (3.11), we get, $\frac{(z-\beta_{1})(z-\beta_{2})}{z^{2}}=\displaystyle\frac{\xi}{b}\overline{q(z)}.$ Hence, we get,
    (3.12) $\frac{\bar{\xi}}{z^{2}}f(z)=\bar{\xi}\frac{\xi}{b}q(z)\overline{q(z)}=\frac{1}{b}\big|q(z)\big|^{2},\text{ for all }z\in\mathbb{T}.$
@@
    We have $\alpha_{1}\alpha_{2}=b~\xi$ and $\alpha_{1}+\alpha_{2}+\frac{1}{\overline{\alpha_{1}}}+\frac{1}{\overline{\alpha_{2}}}=3(1+\xi)$.
    But, $\frac{1}{\overline{\alpha_{1}}}+\frac{1}{\overline{\alpha_{2}}}=\frac{\overline{\alpha_{1}}+\overline{\alpha_{2}}}{\overline{\alpha_{1}}~\overline{\alpha_{2}}}=\frac{\xi}{b}(\overline{\alpha_{1}}+\overline{\alpha_{2}})$.
    Hence, $\alpha_{1}+\alpha_{2}+\frac{\xi}{b}(\overline{\alpha_{1}}+\overline{\alpha_{2}})=3(1+\xi)$.
    Let $\alpha_{1}+\alpha_{2}=r+is$, then we get, b(r+is)+(cosθ+isinθ)(r-is)=3b(1+cosθ+isinθ).
    This gives, $s=a\sin\theta$ and $r=a(1+\cos\theta),$ where $a=\frac{3b}{b+1}.$
@@
    It can be verified that $f(\xi\overline{\alpha_{1}})=0$. We want to prove that $\alpha_{2}=\xi\overline{\alpha_{1}}$. Since, $|\alpha_{1}|,|\alpha_{2}|>1$, it is clear that $\xi\overline{\alpha_{1}}\neq\frac{1}{\overline{\alpha_{1}}}$ and $\xi\overline{\alpha_{1}}\neq\frac{1}{\overline{\alpha_{2}}}$. Hence, it remains to prove that $\xi\overline{\alpha_{1}}\neq\alpha_{1}$.
    Suppose $\xi\overline{\alpha_{1}}=\alpha_{1}$. Let $\arg(\alpha_{i})=\theta_{i}$, for $i=1,2$. We get, $\theta-\theta_{1}=\theta_{1}$ and thus, $\theta=2\theta_{1}$. But, $\alpha_{1}\alpha_{2}=b\xi$ gives $\theta_{1}+\theta_{2}=\theta$. Thus, $\theta_{2}=\theta_{1}$. But, this is not true. Hence, $\xi\overline{\alpha_{1}}\neq\alpha_{1}$.
@@
    The relation between the roots and the coefficients of $f(z)$ gives,
    $\alpha_{1}\alpha_{2}+\alpha_{1}\frac{1}{\overline{\alpha_{1}}}+\alpha_{1}\frac{1}{\overline{\alpha_{2}}}+\alpha_{2}\frac{1}{\overline{\alpha_{1}}}+\alpha_{2}\frac{1}{\overline{\alpha_{2}}}+\frac{1}{\overline{\alpha_{1}}~\overline{\alpha_{2}}}=(8+2\cos\theta)\xi.$
    Using the above relation, it can be seen that $b$ satisfies the polynomial g(x)=x^4-(6+2cosθ)x^3+(4+14cosθ)x^2-(6+2cosθ)x+1. It is easy to observe that for a non-zero real number $\beta$, if $g(\beta)=0$ then $g(\frac{1}{\beta})=0.$
The proof of the main theorem demands different techniques based on the location of the point $\xi=\cos\theta+i\sin\theta$ on the unit circle. We therefore divide the proof into three parts. In fact, the nature of the roots $\alpha_{1},\alpha_{2}$ appearing in Lemma 3.4 when $\cos\theta=0.6$ prompts the authors to make these cases, as can be seen in the following lemma.
@@
Let $\xi=\cos\theta+i\sin\theta\in\mathbb{T}$ $(\xi\neq\pm 1)$. Let $\alpha_{1},\alpha_{2}$ be two complex numbers outside closed unit disc, obtained in part (3) of Lemma 3.4.
@@
    If $\cos\theta>0.6$ then $\arg(\alpha_{1})=\arg(\alpha_{2})$ and $|\alpha_{1}|\neq|\alpha_{2}|.$
@@
    We have $\cos\theta<0.6$. Suppose $\arg(\alpha_{1})=\arg(\alpha_{2})$.
    Now as $\alpha_{1}+\alpha_{2}=a(1+\xi)$ we have $\alpha_{1}=s(1+\xi)$ and $\alpha_{2}=t(1+\xi)$ for some $s,t\in\mathbb{R}$. This gives $\alpha_{1}+\alpha_{2}=(s+t)(1+\xi)=a(1+\xi)$ and $\alpha_{1}\alpha_{2}=st(1+\xi)^{2}=b\xi$.
@@
    Let $D$ be the discriminant of polynomial $h(x)$ then,
    $D=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1+\cos\theta}.$
    Since $\cos\theta<0.6$, we have $\frac{-1}{1+\cos\theta}<\frac{-1}{1.6}$.
    It can be seen that $D<\displaystyle\frac{-b(5b-1)(b-5)}{4(b+1)^{2}}$.
@@
    Recall from part (7) of Lemma 3.4 that $b$ satisfies the polynomial
@@
    Thus there exists $\beta_{0}\in(5,6)$ such that $g(\beta_{0})=0$. We now prove that $g$ has no root in the interval $(1,5)$. Suppose $g$ has a root in the interval $(1,5)$. Let $\beta$ be the smallest such a root of $g.$